OpenStudy (anonymous):

Probability how would you do these problems? A bag contains 3 yellow, 4 blue, 2 red, and 3 orange marbles. You reach in the bag and pick out one marble. What is the probability that you selected an orange marble. An experiment consists of spinning a spinner with eight equal sections numbered 1-8. Find the theoretical probability of spinning a number greater than 6.

OpenStudy (kirbykirby):

1st one: you have 3+4+2+3 = 12 marbles in total There are 3 orange marbles The probability is: \[ \frac{\text{# of ways to get an orange marble}}{\text{total number of marbles}}\] 2nd one: There two ways to get a number greater than 6: namely 7 and 8 The probability is then: \[ \frac{\text{# of ways to get an a number > 6}}{\text{total number of sections}}\]

OpenStudy (anonymous):

4/12 for 1 but that's not an option for answers

OpenStudy (kirbykirby):

there are only 3 orange marbles, not 4 :)

OpenStudy (anonymous):

yeah but im suppose to do 3 divided by 12 correct? the option for answers are a. 1/4 b. 1/12 c. 1/3 d. 1/6

OpenStudy (kirbykirby):

3/12 ... you can simplify the fraction by dividing the top and bottom by 3

OpenStudy (anonymous):

so 1/4?

OpenStudy (kirbykirby):

indeed!

OpenStudy (anonymous):

and the second one would be 3/8?

OpenStudy (anonymous):

no1/4

OpenStudy (kirbykirby):

yup 1/4